Consider the production function Q = 2(KL) 0.5

a)     What is the marginal product of labour and capital

b)     What is the marginal rate of technical substitution of labor for capital

 

c)     What is the elasticity of substitution at a point K = 1, L = 1 if we increase K by one unit?

 


1
Expert's answer
2021-09-03T08:07:35-0400

a) Q = 2(KL)0.5

Q=2K0.5L0.5

Marginal product of labor = Differentiate the above production function with respect to L

dQ/dL=0.5×2K0.5L-0.5

dQ/dL=0.5×2K0.5L-0.5

dQ/dL=K0.5L-0.5

dQ/dL=K0.5/L0.5

Marginal product of capital = Differentiate the above production function with respect to K

dQ/dK=0.5×2K-0.5L0.5

dQ/dL=K-0.5L0.5

dQ/dL=L0.5/K0.5

b) marginal rate of technical substitution of labor for capital =change in capital/chang in labor

="\frac{∆K}{∆L}"

=K0.5L-0.5 /K-0.5L0.5

=K/L

c)  Elasticity of substitution σ is given by

σ=dln(k/l)/dlnMRTS

σ=dln(k/l)/dlnMRTS =1

MRTS=K/L


​σ = 1